Rahim is thinking of retiring within the next 40 years. He believes that he needs at least RM5million in EPF to retire comfortable. He recently bought a terrace house for RM500,000 and he needs to pay mortgages every month for 25 years. Currently, he is putting a side RM4,000 every month until retirement for monthly mortgage payment and retirement. Assuming the nominal annual interest rate including the mortgate rate at 6% but compounded monthly. How much he is still short now to fullfill his ambition to retire with RM5 milion and still manage to pay this monthly mortgage payment?
Answer #1

First we find the PV of amount required at retirement

Using financial calculator
Input: FV=5000000, N=40*12=480 , I/Y=6/12=0.5

Solve for PV as 456310.39

Present value of $4000 kept aside

Using financial calculator
Input: PMT=-4000, I/Y=0.5, N= 480

Solve for PV as 726990.34

PV of Amount short= 456310.39+500000 -726990.34 =

229320.05